OpenStudy (anonymous):

at what average would heat have to be removed from 1.5L of a) water and b) mercury to reduce the liquid's temperature from room temperature to its freezing point in 3 minutes?

OpenStudy (mrnood):

You need to look up a few things first: Density of Water Density of Mercury Specific Heat Capacity of Water Specific Heat Capacity of Mercury Freezing point of Water Freezing point of Mercury. also make an statement of what you will use as 'Room Temperature' Post those here, and I'll help you solve what to do with them.

OpenStudy (anonymous):

okayyy,,,wait,,i will post here

OpenStudy (anonymous):

Density of Water 1000kg/m3 Density of Mercury 13534 kg/m3 Specific Heat Capacity of Water 1 Specific Heat Capacity of Mercury 0.033 Freezing point of Water 0 degrees C Freezing point of Mercury. -38.83 degrees C 'Room Temperature' 20 degrees C

OpenStudy (mrnood):

OK So - you have 1.5 l of water and 1.5 l of Hg (Mercury) Calculate the mass of each of those. Also calculate the change in temperature for each fluid

OpenStudy (anonymous):

15000grams of water and 1500 grams of mercury

OpenStudy (anonymous):

what will i use in change in temperature in both?

OpenStudy (mrnood):

That's correct for water - but Hg is much more dense.... Also - what are the units for specific heat capacity you have posted? It will be easier if you find them in J/gK (or J/g degC)

OpenStudy (mrnood):

Change in temperature is (Room temperature - freezing point)

OpenStudy (anonymous):

specific heAT IS cal/gram C

OpenStudy (mrnood):

So Find the mass of the mercury (you know its volume and density) Find how many calories = 1 Joule

OpenStudy (anonymous):

1 cal is 4.184 joules

OpenStudy (mrnood):

good density = mass/vol Find the mass of the mercury

OpenStudy (anonymous):

1 joule is .2389

OpenStudy (anonymous):

mercury,what it its volume?

OpenStudy (mrnood):

The question tells you - you have a) 1.5 l of water and b) 1.5 l of Mercury

OpenStudy (anonymous):

volume= gram?

OpenStudy (mrnood):

no - you need to know the mass (grams) - and you already know density = 13534 kg/m3 and Volume = 1.5l (You may need to know that 1m^3 = 1000 l) from the above you need to work out mass of 1.5l of mercury

OpenStudy (mrnood):

Let's stick with the water for now OK? What is the mass of 1.5l of water (your answer above was not quite correct I'm afraid)

OpenStudy (anonymous):

mass = density * V = 13600(1.5) = 9066.67, then?

OpenStudy (anonymous):

mass of water = 1000(1.5) = 666.67

OpenStudy (mrnood):

You have 'divided' not multiplied in the above answers - the equation is right but the answer is wrong. Nearly there...

OpenStudy (anonymous):

why divide....? \[\rho= m/V\]

OpenStudy (anonymous):